A toy rocket is launched from a building 148 feet tall at an initial velocity of 222 feet per second.
(a)give the function that describes the height of the rocket in terms of time t.
(b) determine the time at which the rocket reaches its maximum height, and the maximum height in feet
(c) for what time intercal will the rocket be more than 834 feet above ground level?
(d) after how many seconds will it hit the ground?

Answer by stanbon(75887)

A toy rocket is launched from a building 148 feet tall at an initial velocity of 222 feet per second.
(a)give the function that describes the height of the rocket
in terms of time t.
h(t) = -16t^2 + 222t + 148
======================================

(b) determine the time at which the rocket reaches its maximum height,
and the maximum height in feet
Vertex at time = -b/(2a) = -222/(2*-16) = 6.9375 seconds
-----------------------------------------
(c) for what time intercal will the rocket be more than 834 feet
above ground level?
Solve -16t^2+222t+148 = 834
-------------------------------------
(d) after how many seconds will it hit the ground?
Solve -16t^2+222t+148 = 0
=============================
Cheers,
Stan H.